# Towards Solving the Gilbert-Pollak Conjecture via Large Language Models

### v1

@@ E.2 Problem Decomposition and Search Space Reduction
Reduction Strategy for Unbounded $f$. For Cases 2 and 4, where $f>d$, the implicit connection $(D,F)$ (or $(U,F)$) is invalid and thus excluded from the connector graph $t^{*}$. In these scenarios, we apply a strict filtering policy: we only admit splits whose splitting functions are monotonically non-increasing with respect to $f$.
This monotonicity guarantees that the maximum value of the splitting function over the interval $(d,+\infty)$ is bounded by its value at $f=d$. Consequently, we mathematically substitute every occurrence of $f$ with $d$ in the verification functions. This operation effectively reduces the dimensionality of the parameter space $\mathcal{W}$ by one. In these two cases, the variable $f$ is removed entirely, and the verification proceeds on the projection of the hyperrectangles onto the remaining parameters.
  1. 1.
  2. 2.
Appendix F LLM Agent and Prompts
F.1 Prompt for Searching Trapped Regular Point Lemmas
[…]
F.2 Prompt for Searching Valid 4-Point Steiner Tree Lemmas
Prompt for Searching Valid 4-Point Steiner Tree Lemmas PROMPT START Role: You are an expert Computational Geometer and Mathematician. Objective: Derive a simplified, robust condition $P(b,c,d,s,e)$ that guarantees the existence of a specific Steiner Tree. Constraint: The condition you find must be ”Axis-Parallel Segment Closed” (APSC). Tools: You have access to a Mathematica execution environment and a Python execution environment. PHASE 1: Geometric Definitions Construct the following points in the 2D plane based on these vector instructions. A: $(0,0)$ X: From A, move length $1$ at angle $0^{\circ}$. Y: From X, move length $s$ at angle $60^{\circ}$. B: From X, move length $b$ at angle $60^{\circ}$. D: From Y, move length $d$ at angle $0^{\circ}$. P: From Y, move length $c$ at angle $120^{\circ}$. Q: From P, move length $d$ at angle $60^{\circ}$. R: From P, move length $e$ at angle $180^{\circ}$. The domain of the variables is $b\geq 0\;\&\&\;c\geq 0\;\&\&\;d\geq 0\;\&\&\;s\geq 0\;\&\&\;e\geq 0$. Note: Although Points X, Y, P are used for construction, your final task focuses on the Steiner Tree for points {A}, {B}, {C}, {D}. PHASE 2: The Existence Logic (Type {A}{B}-{C}{D}) You are investigating the existence of an ({A}{B})-({C}{D}) type Steiner Tree. For this specific type, define the auxiliary points $U$ and $V$ as follows: 1. Point U: Rotate point {B} $60^{\circ}$ counter-clockwise around {A}. 2. Point V: Rotate point {D} $60^{\circ}$ counter-clockwise around {C}. The Exact Existence Conditions ($E_{exact}$) are: 1. Convexity: The quadrilateral formed by {A, B, C, D} must be convex. 2. Internal Angles: The angles $\angle UCD$, $\angle UDC$, $\angle VAB$, and $\angle VBA$ must all be $\leq 120^{\circ}$. 3. Crossing Angle: Let $\theta$ be the angle between diagonal vectors $\vec{AC}$ and $\vec{BD}$. The condition is $\theta\leq 120^{\circ}$. Relaxation Hint: You may relax this third condition to $\theta\leq 90^{\circ}$ (which is $\vec{AC}\cdot\vec{BD}\geq 0$) if it helps simplify the math. The condition for “convex quadrilateral” can be formulated as follows: For any three adjacent points x, y, z (in clockwise order), the cross product of vectors xy and yz is less than or equal to zero. The condition that the angle between vectors u and v is less than or equal to $120$ degrees can be expressed using the cosine rule or the dot product of vectors, namely: vectors u and v form an angle $\leq 120^{\circ}$ iff $2\,\mathrm{dot}(u,v)+(|u|\,|v|)\geq 0.$ You may relax it to $\mathrm{dot}(u,v)\geq 0$ if possible. PHASE 3: Deriving the Relaxed Condition P The exact conditions above are too complex. You must find a sufficient condition $P(b,c,d,s,e)$ such that $P\implies E_{exact}$. Requirements for Condition P (APSC Property): Definition: A condition $P$ satisfies APSC if, for any line parallel to one of the coordinate axes, the set of points within the domain (variables $\geq 0$) that satisfy the condition forms a single connected interval. To satisfy APSC, $P$ must be the conjunction of inequalities from the following allowed categories only: 1. Linear or Multilinear: Inequalities that are linear in each variable separately. 2. Quadratic: Inequalities in the form $Ax^{2}+Bx+C\geq 0$ (or $\leq 0$), subject to monotonicity rules explained in the prompt. 3. Functional Boundaries: Inequalities in the form $e>f(b,c,s,d)$ or $e<f(b,c,s,d)$, where monotonicity of $f$ must be verified. Coverage Requirement: Your condition $P$ must be general (not hardcoded numbers), but it must strictly cover the following ”Test Box”: {Bottleneck} Important: You MUST verify the APSC Property in the ENTIRE domain, not in the Test Box. PHASE 4: Verification Protocol You must verify your work using Mathematica logic. Step A: Write Mathematica code to compute the exact expressions for the 6 existence inequalities (Convexity + Angles). Step B: Define your relaxed condition condP. Step C: Verify Implication. You must run a command equivalent to this logic (check each inequality separately): Reduce[ !<ExactCondition_i> && condP && b>=0 && c>=0 && d>=0 && s>=0 && e>=0, {vars} ] If this returns False, your condition is valid. If it returns True or an expression, your condition is invalid. Tips: Reformulate $2\,\mathrm{dot}(u,v)+(|u|\,|v|)\geq 0$ into $\mathrm{dot}(u,v)\geq 0\;||\;4\,\mathrm{dot}(u,v)^{2}\leq|u|^{2}|v|^{2}$ for faster speed when verifying. Constraint: Do not put all checks in one Reduce call. It will time out. Check each inequality individually against $P$. Important: You MUST verify the implication in the ENTIRE domain, not in the Test Box. PHASE 5: Output Requirements 1. Detailed Derivation: Show how you calculated coordinates and how you derived the relaxed inequalities. 2. APSC Proof: Provide a logical argument explaining why your condition $P$ satisfies the APSC property, specifically addressing the Linearity/Monotonicity requirements for each part of your condition. 3. Mathematica Output: Paste the code and the output confirming the False result for the implication checks. 4. C++ Implementation: Provide a C++ code block containing these exact functions: // Returns true if your condition \$P\$ is met. bool steiner_cond(double b, double c, double d, double s, double e) { // Your condition here } // Returns the total length of the ({A}{B}-{C}{D}) type Steiner tree, which is exactly |UV|, the distance between U and V. double steiner_length(double b, double c, double d, double s, double e) { // Return the distance between U and V } Important: Do NOT use the Test Box values inside the C++ code. The code must be a general algebraic formula. If you fail to find a valid condition after rigorous attempts, return false and $0.$ — END OF PROMPT —

### v2

@@ H.5 Verification Procedure
Reduction Strategy for Unbounded $f$. For Cases 2 and 4, where $f>d$, the implicit connection $(D,F)$ (or $(U,F)$ and $(V,F)$) is invalid and thus excluded from the graph $t^{*}$. In these scenarios, we apply a strict filtering policy: we only admit splits whose splitting functions are monotonically non-increasing with respect to $f$.
Hence, in the region $f\geq d$, it suffices to verify the boundary $f=d$. For each function $F$, define $F^{\partial}$ as the reduced function obtained from $F$ by setting $f=d$ and removing $f$ from the input variables. The verification target is thus reduced to
$\max_{{\bm{w}}\in\mathcal{W}^{\partial}}\min_{F\in\mathcal{F}}F^{\partial}({\bm{w}})\leq 0,$
where $\mathcal{W}^{\partial}$ denotes the reduced parameter space obtained by removing the $f$-coordinate from $\mathcal{W}$, so that $\mathcal{W}^{\partial}=[0,+\infty)^{n-1}$.
Note that $F^{\partial}$ preserves both convexity and coordinatewise non-increasingness of $F$. In particular, its non-increasingness with respect to $d$ follows because increasing $d$ simultaneously increases the original $d$-entry and the substituted $f$-entry, and $F$ is non-increasing in both variables.
It remains to handle the four cases above separately. For each case, we construct the corresponding verification functions using Lemma 2736, with the boundary-reduced forms applied in Cases 2 and 4, and then verify the resulting inequalities via Algorithm 1. This completes the verification procedure.
The complete implementation for generating and verifying the certificates is available at https://github.com/keyisi2006/Steiner-Ratio/tree/main/certificate, allowing the computer-assisted verification to be checked independently.